Presto Company makes radios that sell for $28 each. For the coming year, management expects fixed costs to total $278,880 and variable costs to be $16.80 per unit.

(a) Compute the break-even point in dollars using the contribution margin (CM) ratio.

Break-even point

$


(b) Compute the margin of safety ratio assuming actual sales are $944,715. (Round margin of safety ratio to 2 decimal places, e.g. 10.50%.)

Margin of safety

%


(c) Compute the sales dollars required to earn net income of $85,120.

Required sales

$
Answer #1

a)the break-even point in dollars using the contribution margin:

Breakeven points = fixed cost/(selling price per unit-variable cost per unit)*selling price per unit

=(278880/(28-16.80))*28

=$697200

------------------------

b. margin of safety ratio assuming actual sales are $944,715.

MOS ratio = (Budgeted Sales ? Break-even Sales)/Budgeted Sales

=(944715-697200)/944715

=0.26

------------------------------

c. sales dollars required to earn net income of $85,120:

Let the number of units required be x

=>(x*28)-(x*16.8)-278880 = 85120

=> x= (85120+278880)/11.2

=> x = 32500 units

Sales required = 32500*28

=$910000
Answer #2

Answer:-

sales price (a) $ 28
variable price (b) $ 16.8
contribution (a-b) $ 11.2
fixed expenses $ 2,78,880

A) break even point = fixed expenses / (selling price - variable price)

Break even point = 2,78,880 / (28 - 16.8)

break even point = $ 24,900

Answer (B) :-

margin of safety = actual sales - break even sales

actual sales = $ 9,44,715 (given in ques)

break even sale = $ 24,900 (compuate from answer A)

margin of safety = 9,44,715 - 24,900

= $ 9,19,815

Answe (C) :-

net income = $ 85,120

calculation of sales volume to earn desried profit =

sales in $ = (fixed expenses + profit) * selling price / selling price - variable cost

= (278880 + 85120) * 28 / (28 - 16.8)

= 364000 * 20 / 11.2

sales in $ = 9,10,000